# General Place Recognition Survey: Towards Real-World Autonomy

### v1

@@ VI-A Public Datasets
Table II provides a summary of several commonly utilized PR datasets, emphasizing key factors such as scenarios, scale, sensor types, and the associated challenges.
@@ VI-A1 VPR Datasets
Related datasets predominantly cater to various environmental conditions, including illumination [107] and seasons [169]. The Nordland dataset [169], captured using a train-mounted camera, is particularly notable for benchmarking condition-invariant or large-scale PR tasks. In the realm of lifelong PR, Warburg et al. [16] introduced the most extensive VPR dataset to date, covering urban and suburban settings over a span of seven years and documenting various condition changes. The ALIO dataset [173] presents a comprehensive dataset for the VTRN task, including raw aerial visuals and corresponding satellite imagery. Furthermore, Project Aria [175] utilizes AR glasses to provide a dataset of $275$ sequences over $15$ months from a construction site, which is crucial for exploring lifelong PR with structural and appearance changes.
@@ VI-A2 LPR Datasets
The datasets discussed herein extensively overlap with those for VPR, with a primary focus on urban environments. Notable autonomous driving datasets, including KITTI [170] and Oxford RobotCar [171], are particularly valuable for evaluating PR in diverse open-street contexts. The Newer College dataset [176] is a campus-scene dataset that features LiDAR and stereo-inertial sensors, which is well-suited for VPR, PR, and cross-modal PR evaluations. Additionally, the ALITA dataset presents large-scale, LiDAR-centric data encompassing $50$ city-scale trajectories with $120$ overlaps and $80$ campus-scale trajectories with $150$ overlaps. This dataset is versatile, supporting a broad spectrum of algorithm evaluations, including large-scale and cross-domain PR, multi-agent map merging, and lifelong learning.
@@ VI-A3 RPR Datasets
RPR datasets typically feature extreme environments under various weather conditions, including foggy and snowy days, where Radar technology demonstrates significant advantages. Key datasets such as the Oxford RoboCar Radar [174] and Boreas [15] showcase Radar’s unique capabilities in challenging visibility conditions.
@@
Figure 12: Star-diagram for PR Evaluation on two datasets. We compare five PR methods in terms of four properties by reporting average precision (AP), number of total and trainable parameters ($N_{\text{total}}$ and $N_{\text{train}}$, and average computation time for inference at each frame (CT). “AP day-day” indicates that query and database images are both captured during daytime.
@@ VI-B New Perspective of Evaluation
  • Training and Inference Cost: Analyzes computational demands, including NP and CT required by a PR algorithm.
Fig. 12 illustrates two star diagrams comparing the properties of five SOTA PR methods (without fine-tuning) using two subsets of the ALITA-campus dataset [126]. We created the Dataset $1$ by selecting two sequences from the same location captured during daytime and nighttime, respectively, to assess the condition-invariant property and generalization ability of the PR methods. The notation “day-night” indicates the test setting that daytime images serve as the database, while nighttime images form the query set. Unless specified, database and query images are chosen randomly. The Dataset $2$ follows a similar structure but focuses on evaluating the viewpoint-invariant property. To aid researchers in developing novel PR methods, we provide a standardized implementation and evaluation framework detailed at GPRS-core [^1] https://github.com/MetaSLAM/GPRS_core, which will be further explained in the subsequent section.
VI-C Supported Libraries
Recent advancements have led to the development of several comprehensive libraries aimed at enhancing long-term PR systems. OpenSeqSLAM2.0 [177] enhances sequence matching by providing a detailed analysis of SeqSLAM’s key components. VPR-Bench [6] integrates $12$ datasets and $10$ VPR methods, complete with evaluation metrics to benchmark new PR techniques. Additionally, Kapture [178] is an open-source library that supports visual localization and structure-from-motion, including implementations of VPR.
We publicly release the GPRS-core, which is a versatile framework taliored for PR benchmarking and support multiple input modalities such as point clouds and images. The usage of GPRS-core involves three main steps: (1) training a PR model and saving the weights; (2) applying these weights across multiple datasets; and (3) evaluating model performance with diverse metrics for benchmark comparisons. Notably, GPRS-core reimplements over $11$ SOTA methods: (1) VPR Methods: NetVLAD [25], CosPlace [28], MixVPR [82], SVTNet [93], and TransVPR [29]. (2) LPR Methods: BEVPlace [55], LoGG3D [52], MinkLoc3D [50], MinkLoc3Dv2 [179], and PointNetVLAD [45]. (3) Fusion-Based Method: MinkLoc++ [35]. We invite the academic community to contribute to enhancing the package, thereby advancing PR research.
@@ VII Conclusion
The escalating complexity of mobile robots necessitates the development of lifelong navigation systems capable of autonomous, long-term operations in large-scale environments. PR, which enables robots to identify previously visited locations under appearance changes and viewpoint differences, has emerged as an essential technology for robotic autonomy. This paper has charted significant advancements in PR over recent decades, including the discussion of the definition of PR and typical methods on representations, strategies against challenges, and applications. We start providing a clear problem formulation that is closed to the need of robotic navigation of position-based PR. This sidesteps the ambiguities associated with overlap-based PR definitions which are aligned with image retrieval tasks.
Addressing the critical question of “representing a place”, we explore the evolution from handcrafted feature extraction to data-driven methodologies. Methods of place representation is benefiting from research in other fields, particularly the great strides being achieved in computer vision and nature language processing in the fields of novel neural network architectures, open-set object detection, semantic segmentation, and so on. Developments in high-level place representations (i.e., graphs and implicit embedding) have collectively fostered the paradigm shift. This shift not only simplifies the PR challenges but also enhances model generalization abilities and provides more opportunities to define PR architectures.
The challenge of accurately “recognizing the correct place” for real-world robots introduces five fundamental challenges: appearance change, viewpoint difference, model generalization, efficiency on resource-constrained platforms, and uncertainty estimation of the methods’ output. Key solutions are reviewed, while unsolved issues still exist. The trend shows that the development of methods from the precision race in open-datasets to solution of real-world problems encountered.
In recent years, the development of PR has advanced in parallel with SLAM algorithms. Fortunately, an increasing number of studies are now proposing or incorporating SOTA PR methods to enhance navigation systems. Looking towards real-world navigation, we question ”how to bridge the gap between PR algorithms and applications?” Highlighting PR’s vast potential, we showcase applications from large-scale navigation and visual terrain navigation to multi-agent systems and beyond, including VR/AR and crowd-sourced mapping. This field presents numerous opportunities for exploration. Additionally, the indispensable contributions of PR datasets, evaluation metrics, and open-source libraries in expediting the advancement of PR are discussed.
In conclusion, we underscore the immense potential of PR in shaping the future of robotic navigation and autonomy. By inviting researchers to engage with our open-source projects, we aim to collectively advance towards the realization of generalized PR, a goal that holds promise for transforming both robotic systems and their applications in our world.

### v2

@@ VI-A Public Datasets
Table II provides a summary of several commonly utilized PR datasets and highlights the key factors.
@@ VI-A1 VPR Datasets
Related datasets predominantly cater to various environmental conditions, including repetitive structures [207], illumination [93, 208], and seasons [202]. The $24$/$7$ Tokyo [93] and Pitts30303030k [207] are two classical VPR datasets for their features in street-view imagery. The Nordland [202], SVOX [208], and Boreas datasets [14] are designed for cross-seasonal VPR, with the former covering natural environments and the latter two focusing on urban settings. The NYC-Event-VPR [209] dataset was proposed for event-based VPR in dynamic urban environments. In the realm of lifelong PR, Warburg et al. [15] introduced the most extensive VPR dataset to date, covering urban and suburban settings over a span of seven years and documenting various condition changes. The ALIO dataset [205] presents a comprehensive dataset for the VTRN task, including raw aerial visuals and corresponding satellite imagery.
@@ VI-A2 LPR Datasets
LPR shares many benchmark datasets with VPR. Urban driving datasets like KITTI [210] and Oxford RobotCar [203] are valuable for evaluating PR in open-road scenarios. Campus environments are represented by the Newer College dataset [211], which provides synchronized LiDAR and stereo-inertial data. The ALITA dataset [157] expands evaluation versatility with $50$ city-scale ($120$ overlapping pairs) and $80$ campus-scale trajectories ($150$ overlapping pairs), supporting tasks ranging from cross-domain recognition to lifelong SLAM. Emerging natural environment datasets [212, 213, 214] address challenges like structural ambiguity and dynamic vegetation, driven by growing applications in forestry, agriculture, and subterranean robotics.
@@ VI-A3 RPR Datasets
RPR datasets typically feature extreme environments under various weather conditions, including foggy and snowy days, where Radar technology demonstrates significant advantages. Key datasets such as the Oxford RoboCar Radar [206], MulRan [215], and Boreas [14] showcase Radar’s unique capabilities in challenging visibility conditions.
@@
Figure 12: Star-diagram for PR evaluation on two datasets. We compare five PR methods in terms of four properties by reporting average precision (AP), number of total and trainable parameters ($N_{\text{total}}$, $N_{\text{train}}$), and average computation time for inference at each frame (CT). AP day-day indicates that query and database images are both captured during daytime.
@@ VI-B New Perspective of Evaluation
  • Training and Inference Cost: Analyzes computational demands, including NP and CT required by a PR algorithm. (extract descriptors)
Fig. 12 presents star diagrams comparing five SOTA PR methods (without fine-tuning) using two subsets of the ALITA-campus dataset [157]. Dataset 1 comprises two sequences from the same location captured during daytime and nighttime, was designed to evaluate condition invariance and generalization. The day-night notation denotes a test setting where daytime images form the database and nighttime images serve as queries. Unless stated otherwise, database and query images are randomly selected. Dataset 2 adopts a similar structure but focuses on viewpoint invariance. For fair comparison and visualization of method performance, we provide evaluation scripts in this repository [^1] https://github.com/MetaSLAM/GPRS
@@ VII Conclusion
The increasing sophistication of mobile robots necessitates navigation systems that sustain indefinite autonomy in expansive, dynamic environments. PR, which enables robots to identify previously visited locations despite appearance changes and viewpoint variations, has become crucial for robotic autonomy. This survey outlines significant advancements in PR, detailing its definition, typical representations, strategies for overcoming challenges, and diverse applications. We present a formulation for effective PR, associating it with the need for robotic navigation.
Focusing on the core challenge of “representing a place,” we examine the shift from handcrafted features to data-driven methods, benefitting from progress in computer vision and maching learning, especially in neural networks, open-set object detection, and semantic segmentation. This paradigm shift in high-level representations simplifies PR challenges, enhances model generalization, and creates new opportunities for PR architecture design. The real-world deployment of PR faces five primary challenges: appearance changes, viewpoint variations, model generalization, resource efficiency, and output uncertainty estimation. We review key solutions to these challenges, highlighting the research community’s gradual pivot from dataset-driven methodologies to systems validated in real-world environments. PR development has paralleled SLAM advancements, with an increasing number of studies integrating SoTA PR methods to improve navigation systems. Real-world applications now stretch from large-scale and visual terrain navigation to multi-agent systems, VR/AR, and crowdsourced mapping. The contributions of PR datasets, evaluation metrics, and open-source libraries have been instrumental in advancing the field.
In conclusion, PR holds immense potential for advancing robotic autonomy. Through this paper and our future efforts, we aim to accelerate progress toward generalized PR, shaping the future of robotic systems and their applications.
VIII Acknowledgement
The authors thank Jingwen Yu, Jianxing Shi, Xinyi Chen, and Shuyang Zhang for their constructive feedback, and Profs. Michael Milford and Dimitrios Kanoulas for insights on PR definitions and applications. We also acknowledge ChatGPT (OpenAI) and DeepSeek (DeepSeek-AI) for text refinement.